OpenStudy (anonymous):

hey someone please help me with this.please. A light at eye level stands 20 ft. from a house and 15 ft. from the path leading from the house to the street.A man walks along the path at 6 ft per sec. How fast does his shadow move along the wall when he is 5 ft. from the house?

OpenStudy (anonymous):

time-rates

OpenStudy (anonymous):

you got a picture?

OpenStudy (anonymous):

no i don't have.

OpenStudy (anonymous):

the ratio of distances to the man and to the house gives you the ratio for the rates: 20/15 = x/6

OpenStudy (anonymous):

i can't see it. the light is not on the path correct?

OpenStudy (anonymous):

yes based on the problem it's not bec. its 15 ft. from the path

OpenStudy (anonymous):

hey tonks do you know how to solve it?

OpenStudy (anonymous):

yes 20/15 = x/6 reduce 4/3 = x/6 mult by 6 24/3 = x 8 = x

OpenStudy (anonymous):

tonks i'm sorry the connection is freaking out..

OpenStudy (anonymous):

@ tonks can you show the details..pls. solve it step by step please.

OpenStudy (anonymous):

just really need your help please...

OpenStudy (anonymous):

i must be really dumb because i cannot even draw this picture. the man is walking along the path to the house, but where is the light?